# Generation of Random (Generalized) Orthogonal Matrices

### v1

@@ 3.1   Case 1: $S$ is Symmetric
Thus, generating an element $A$ of $O_{S}(N)\cap O(N)$ is equivalent to generating a block diagonal orthogonal matrix $B$. Since the block matrices $B_{i}$ are clearly independent due to the block matrix multiplication and the orthogonality definition $B^{T}B=I_{n}$, one can generate $A$ by the following steps:
@@
    For each $i$, extract a random $k_{i}\times k_{i}$ orthogonal matrix $B_{i}$ by applying the QR decomposition to $A_{i}$.
@@
    To ensure that the matrices $B_{i}$ (and hence $B$) are suitably random, apply the QR decomposition in such a way that it is unique. Specifically, enforce the diagonal entries of the upper triangular matrices $R_{i}$ obtained from the QR decomposition of $A_{i}$ to be positive (see [3]).
@@
    Form the block diagonal matrix $B=\text{diag}(B_{1},B_{2},\ldots,B_{m})$.
  5. 5.
@@
2:Output: An $S$-orthogonal matrix $A$
@@ 3.2   Case 2: $S$ is Skew-Symmetric
The numbers $\lambda_{k_{j}}$ are the eigenvalues of $T$ with positive imaginary part, as described in Theorem 3.
@@
We next provide a characterization of the class of orthogonal matrices satisfying equation (16). In other words, we characterize the groups $O_{J}(2N)\cap O(2N,\mathbb{R})$, where $J$ has the above form. The characterization is summarized in the following proposition that generalizes a well-known fact about symplectic matrices (see [[5], p. 16]).
@@
where the last equality follows from the conditions in the first observation.
@@
hence $Q\in O(2n,\mathbb{R})$. Moreover, it follows by the third observation that $Q\in O_{J}(2N)$, which proves the inclusion $U(n)\subset Sp(n)\cap O(2n,\mathbb{R})$.
Conversely, suppose that $Q\in O_{J}(2N)\cap O(2n,\mathbb{R})$. Then
@@
which implies that $Q\in U(n)$, so $O_{J}(2N)\cap O(2n,\mathbb{R})\subset U(n)$. $\blacksquare$
@@
Thus, similar to case 1, since the block matrices $B_{i}$ are clearly independent, one can apply the QR factorization to random invertible complex matrices as described in Algorithm $1$ to generate the desired matrix (see Algorithm $2$ on page $12$).
@@
2:Output: An $S$-orthogonal matrix $A$

### v2

@@ Abstract
Keywords: Bilinear Forms, Orthogonal Group, Unitary Group, Symplectic Group, Indefinite Orthogonal Group, Schur Decomposition, QR Factorization, and Polar Decomposition.
@@ 3.1   Case 1: $S$ is Symmetric
Thus, generating an element $A$ of $O_{S}(N)\cap O(N)$ is equivalent to generating a block diagonal orthogonal matrix $B$. Since for all $i\neq j$, every entry of a block matrix $B_{i}$ is pairwise independent and identically distributed with the entries of a block matrix $B_{j}$ due to the block matrix multiplication and the orthogonality definition $B^{T}B=I_{n}$, one can generate $A$ by the following steps:
@@
    For each $i$, extract a random $k_{i}\times k_{i}$ orthogonal matrix $B_{i}$ by applying the QR decomposition to $A_{i}$. To ensure that the matrices $B_{i}$ (and hence $B$) are suitably random, apply the QR decomposition in such a way that it is unique. Specifically, enforce the diagonal entries of the upper triangular matrices $R_{i}$ obtained from the QR decomposition of $A_{i}$ to be positive (see [3]). An alternative approach to generating the random orthogonal matrices $B_{i}$ is to employ the polar decomposition. When applied to an invertible matrix, the polar decomposition is unique, thereby ensuring that the orthogonal matrices are appropriately generated.
@@
    Form the block diagonal matrix $B=\text{diag}(B_{1},B_{2},\ldots,B_{m})$.
@@
2:Output: An orthogonal $S$-orthogonal matrix $A$
@@ 3.2   Case 2: $S$ is Skew-Symmetric
The numbers $\lambda_{k_{j}}$ are the eigenvalues of $T$ with positive imaginary parts,  as described  in Theorem 3.
@@
We next provide a characterization of the class of orthogonal matrices satisfying equation (16). In other words, we characterize the groups $O_{J}(2N)\cap O(2N)$, where $J$ has the above form. The characterization is summarized in the following proposition that generalizes a well-known fact about symplectic matrices (see [[5], p. 16]).
@@
where the last equality follows from the conditions in the first and second observations.
@@
hence $Q\in O(2n)$. Moreover, it follows by the third observation that $Q\in O_{J}(2N)$, which proves the inclusion $U(n)\subset Sp(n)\cap O(2n)$.
Conversely, suppose that $Q\in O_{J}(2N)\cap O(2n)$. Then
@@
which implies that $Q\in U(n)$, so $O_{J}(2N)\cap O(2n)\subset U(n)$. $\blacksquare$
@@
Thus, similar to case 1 in section 3.2, since for all $i\neq j$, every entry of a block matrix $B_{i}$ is pairwise independent and identically distributed with the entries of a block matrix $B_{j}$, one can apply the QR factorization to random invertible complex matrices as described in Algorithm $1$ to generate the desired matrix (see Algorithm $2$ on page $12$).
@@
2:Output: An orthogonal $S$-orthogonal matrix $A$